Chords, radius and diameter
Given: chord lenght=16
a line from the center to the chord=6
What is the formula to find the diameter?
Thanks...Kayley

Answer by scott8148(6628)

if the line is from the center of the circle to the center of the chord...it is a perpendicular bisector

the line, half the chord, and the radius form a right triangle; with the radius as the hypotenuse

so 6^2+(16/2)^2=r^2 __ 36+64=r^2 __ 10=r

if the radius is 10, then the diameter is 20